OpenStudy (anonymous):

Can anyone help me with this question? -Given the measures of 3 angles of a triangle, solve for x. x+10, x-20, x+25

OpenStudy (anonymous):

Those are the measurements of the three angles?

OpenStudy (tyteen4a03):

Remember the angle sum of triangle formula: \(\angle A + \angle B + \angle C = 180\)

OpenStudy (raden):

hint : the sum of interior angles of a triangle is 180 degrees so, x+10+x-20+x+25 now, solve for x

OpenStudy (anonymous):

Yeah, I know the triangle formula. I just don't understand how to do this question at all. Like,what is the first step?

OpenStudy (anonymous):

Theres not 3 triangles. Your completely off lol @aylin

OpenStudy (tyteen4a03):

@barnoldwrl He said "angles".

OpenStudy (anonymous):

@tyteen4a03 it's actually she. :P

OpenStudy (anonymous):

Whooops. little tired lol! sorry @aylin

OpenStudy (tyteen4a03):

@Aylin Noted. @barnoldwrl Look at @RadEn 's explanation and see if you get the hang of it. Basically he's saying all these 3 angles will add up to be 180 degrees, so you can use this fact to build an equation and solve for x.

OpenStudy (anonymous):

I need to know how to build an equation in order to do so!! :/

OpenStudy (tyteen4a03):

@barnoldwrl You are given that the three angles added up equals 180. The three angles added up are x+10+x-20+x+25, and they equal 180, so just put them together and get x+10+x-20+x+25 = 180.

OpenStudy (anonymous):

Okay i understand that, but how do i solve for x? this is so confusing..lol well for me atleast! I suck so bad in geometry its not even funny!

OpenStudy (tyteen4a03):

It's actually algebra. ;) Firstly group the like terms. Basically this means variables go with variables, and constants (numbers without variables) go with constants. After some grouping we have \(x+x+x+10-20+25 = 180\). Now simplify the terms and you should get a better equation to solve.

OpenStudy (anonymous):

Im so still lost. lol

OpenStudy (tyteen4a03):

Which part are you stuck on?

OpenStudy (anonymous):

Just grouping the like terms. I just dont know how to do it. Im so behind.

OpenStudy (tyteen4a03):

Grouping variables work like this: if you have x and another x, add it up and it becomes 2x. Simple arithmetic but with a variables after the result. Similarly 2x + 4x = 6x.

OpenStudy (anonymous):

Yeah i get that.

OpenStudy (anonymous):

So, theres 3x right? since there is 3 x's?

OpenStudy (tyteen4a03):

Correct.

OpenStudy (anonymous):

SO whats the equation now?

OpenStudy (tyteen4a03):

You just done grouping the variable terms, so now you have 3x+10-20+25 = 180.

OpenStudy (anonymous):

Whats the next step?

OpenStudy (anonymous):

@tyteen4a03

OpenStudy (tyteen4a03):

Simplify the 10-20+25 bit and continue solving.

OpenStudy (anonymous):

15?

OpenStudy (tyteen4a03):

Correct. Now that you have 3x + 15 = 180, you can solve this equation.

OpenStudy (anonymous):

is it 30?

OpenStudy (anonymous):

55!? @tyteen4a03

OpenStudy (tyteen4a03):

Correct.

OpenStudy (anonymous):

Can you help with a few more or no?

OpenStudy (anonymous):

@tyteen4a03

OpenStudy (tyteen4a03):

Open a new question for these.

OpenStudy (anonymous):

Okay! I am now.